OpenStudy (anonymous):

Use the Change of Base Formula and a calculator to evaluate the logarithm, correct to six decimal places. Use either natural or common logarithms. log(base7) 10

OpenStudy (anonymous):

This is a good one because I'm not sure how to select a specific base in my sci. cal. ha ha

OpenStudy (anonymous):

\[\frac{\ln(10)}{\ln(7)}\]

OpenStudy (anonymous):

you could use: \[\log_ba = \frac{\log(a)}{\log(b)}\] Then you would have: \[\log_7(10) = \frac{\log(10)}{\log(7)} = \frac{1}{\frac{\log(7)}{\log(10)}} = \frac{1}{\log7}\]

OpenStudy (anonymous):

sweet!... I forgot about that